lolokrat:

Write the equation of the line passing through (3, 7) that is perpendicular to 5x − 6y = 4.

lolokrat:

i got \[y=\frac{ 5 }{ 6 }x - \frac{ 27 }{ 6 }\] but it is wrong

lolokrat:

here were my steps: to find m: 5x-6y=4 -6y=-5x+4 y=-5/6x - 4/6 m=5/6 line through (3,7) with slope m=5/6 y-y1=m(x-x1) y-7=5/6(x-3) y+7=5/6x-15/6 y=5/6x-15/6-42/6 y=5/6x-27/6

lolokrat:

here is the example problem that i referenced

surjithayer:

eq. of any line pependicular to ax+by=c is bx-ay=d where d is a constant. eq. of liine perpendicular to 5x-6y=4 is 6x+5y=c \[\because~it ~passes~ through~ (3,7)\]6(3)+5(7)=c c=18+35=53 so reqd. line is 6x+5y=53

lolokrat:

ah thank you surj